OpenStudy (anonymous):

5(4x+5)+4(6x+5)+2=44x+48

OpenStudy (anonymous):

solve for x.

OpenStudy (anonymous):

w8

OpenStudy (anonymous):

20x+25+24x+20= 44x+48 44x+45= 44x+48 45 not equal to 48 as you can see this equality doesn't hold ...so no value of x (ans.) or either you hav posted wrong question.

OpenStudy (anonymous):

also i just can't get this: \[\frac{ 1 }{ 5 }x+ \frac{ 1 }{ 3 }=1(\frac{ 2 }{ 3 }x+2)\]

OpenStudy (anonymous):

regarding the answer above ! thank you sir! answer no value :'D

OpenStudy (anonymous):